Grade 11Mechanics

You throw a ball straight up with an initial velocity of 15m/s.it passes a tree branchon the way up at height of seven .how much addition al time will pass before the ball passese the tree branch on their ay back down

Aastha Shekhawat
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
u=15 m/s 
total height reached by ball be H 
KEi is initial kinetic energy 

velocity at 7 m height be v 
at a height of 7 m it contains both KE &PE 
KEi=PE+KE 
0.5mu^2=mgh+0.5mv^2 [h=7 m] 
v=9.4 m/s 

at top height velocity is 0 let the velocity at top be vf 
we know 
v=u+at 
vf=v-gt [a=-g since the motion is in upward direction] 
t=0.94 s 

so total time to come back to the branch=2*t=2*0.94=1.88 s ~1.91 s